# Complete Security and Privacy for AI Inference in Decentralized Systems

### v1

@@ 4.2 Nesa’s Innovation of NTEE
By following these steps, Nesa ensures that the TEEs needing secure communication can transmit data securely and efficiently. This method reduces the reliance on the orchestrator for data routing, minimizes latency, and leverages the computational strengths of both GPU TEEs and CPU TEEs. The secure communication channels maintain data privacy and integrity throughout the decentralized AI inference process.
Innovation 2: Heterogeneous TEE Scheduling. In Nesa’s system, TEEs are heterogeneous, meaning they vary in type and computational power, with some nodes supporting GPU TEEs and others only CPU TEEs. To manage this diversity, Nesa employs a dynamic scheduling strategy based on initial attestation results. This attestation determines whether a node supports GPU TEE or CPU TEE only. The motivation behind this is to leverage the strengths of each type of TEE, ensuring that computationally intensive tasks are handled by more capable nodes, thereby optimizing overall system performance.
@@
Innovative Motivations. The innovations in Nesa’s use of TEEs are driven by several key motivations:
  • Security and Privacy: Ensuring the confidentiality and integrity of user data and AI models is paramount. By using TEEs and secure communication channels, Nesa guarantees that sensitive information is protected throughout the computation process.
@@
  • Scalability: As the number of nodes increases, the system must efficiently manage communication and computation across a distributed network. Nesa’s innovations enable seamless scaling without compromising on performance or security.
By implementing these innovations, Nesa effectively manages the decentralized AI inference process, ensuring secure, efficient, and optimal performance across heterogeneous TEEs. This approach leverages the strengths of both GPU and CPU TEEs, providing robust security measures and dynamic resource allocation to meet diverse computational needs.
5 Conclusions and Future Directions
Artificial intelligence and machine learning have made remarkable advancements, especially with the development of large language models like ChatGPT and diffusion models such as DALLE-3 and Sora. These foundation models have demonstrated impressive capabilities, but they also raise significant security and privacy concerns, particularly in distributed and decentralized computing scenarios. Ensuring the integrity of inference results and maintaining data privacy are important challenges. Nesa addresses these challenges by offering adaptive security and privacy solutions tailored to different scenarios. For critical inference tasks, where high accuracy and confidentiality are essential, we recommend using a zero-knowledge Decentralized Proof System (zkDPS) for model verification and Sequential Homomorphic Encryption (SHE) for data privacy. For general inference tasks, where efficiency and speed are prioritized, we propose Consensus-based Verification (CBV) for model integrity and leverage Split Learning (SL) for data encryption. Additionally, Nesa’s Trusted Execution Environment (NTEE) offers a hardware-based approach that can be used as an orthogonal solution, addressing both verification and encryption needs simultaneously. This hybrid strategy allows us to balance security, privacy, and efficiency, providing users with the flexibility to choose the most suitable solution for their specific needs.
Looking ahead, we identify several key areas for further research and development to enhance our security and privacy frameworks. First, we aim to optimize zkDPS to reduce computational overhead and increase processing speed, making it more feasible for real-time applications. Second, recognizing that different AI models may require tailored zero-knowledge proof designs, we plan to develop ZK templates that can be easily adapted to various models, ensuring robust security across diverse AI applications. Lastly, given the multiple security and privacy options available, we plan to build an automated framework that can recommend and choose the best set of approaches based on specific user needs and application scenarios, which can be built on top of meta-learning and/or fine-tuned LLMs. This framework will enhance user experience by providing customized, optimal security solutions. By pursuing these future directions, we aim to continually improve the robustness, efficiency, and adaptability of our security and privacy solutions, maintaining Nesa’s leadership in secure and private AI deployment.

### v2

@@ C.2 Our Innovations in Specialized TEE
By following these steps, we ensure that the TEEs needing secure communication can transmit data securely and efficiently. This method reduces the reliance on the orchestrator for data routing, minimizes latency, and leverages the computational strengths of both GPU TEEs and CPU TEEs. The secure communication channels maintain data privacy and integrity throughout the decentralized AI inference process.
Innovation 2: Heterogeneous TEE Scheduling. In our system, TEEs are heterogeneous, meaning they vary in type and computational power, with some nodes supporting GPU TEEs and others only CPU TEEs. To manage this diversity, we employ a dynamic scheduling strategy based on initial attestation results. This attestation determines whether a node supports GPU TEE or CPU TEE only. The motivation behind this is to leverage the strengths of each type of TEE, ensuring that computationally intensive tasks are handled by more capable nodes, thereby optimizing overall system performance.
@@
Innovative Motivations. The innovations in our use of TEEs are driven by several key motivations:
  • Security and Privacy: Ensuring the confidentiality and integrity of user data and AI models is paramount. By using TEEs and secure communication channels, we guarantee that sensitive information is protected throughout the computation process.
@@
  • Scalability: As the number of nodes increases, the system must efficiently manage communication and computation across a distributed network. Our innovations enable seamless scaling without compromising on performance or security.
By implementing these innovations, we effectively manage the decentralized AI inference process, ensuring secure, efficient, and optimal performance across heterogeneous TEEs. This approach leverages the strengths of both GPU and CPU TEEs, providing robust security measures and dynamic resource allocation to meet diverse computational needs.